# EvoC2Rust: A Skeleton-guided Framework for Project-Level C-to-Rust Translation

### v3

@@ 4.4. Scalability on Larger Projects (RQ4)
To evaluate the scalability and performance of our method on large-scale codebases, we conduct experiments on 10 larger projects (9,973 - 91,588 LoC) from RepoTransBench (repotransbench), a benchmark derived from highly-starred GitHub repositories. During module translation, we exclude functions that depend on third-party libraries, as our current implementation primarily supports dependencies on the ISO C standard library. Using DeepSeek-V3 as the backbone model, EvoC2Rust achieves acceptable performance, with an ICompRate of 75.61% and a SafeRate of 96.20%. We do not report line acceptance rates due to the prohibitive cost of manually creating verified Rust references for projects at this scale.
@@
Figure 3. Translation performance at different project scales.
Figure 4. Translation time at different project scales.
@@ 4.5. Case Study
Figure 5. A case study on translating the rb_tree_rotate function.
@@ 5. Limitations and Threats to Validity
Internal Validity. The main internal threat comes from correctness verification, which currently relies on predefined test cases. Adopting more robust validation techniques like fuzzing (KadronNPBPS24) and self-debugging (KangCYL25; shi2024code) could enhance test coverage and uncover subtle semantic errors. For line acceptance rate measurement, we use Claude-generated outputs verified by human experts as ground truth. While this manual validation ensures correctness beyond automated testing, it may not scale to larger datasets. Future work should explore semi-automated verification to balance accuracy and efficiency.
@@ 6.1. C-to-Rust Translation
Rule-based Methods rely on manually designed rules to transform C code into Rust. The C2Rust transpiler (c2rust) exemplifies this approach by using Clang’s AST to generate Rust code, though the output relies heavily on unsafe blocks. Subsequent work has aimed to improve safety and idiomaticity: Emre et al. (EmreSDH21; EmreSDH23) used compiler feedback for pointer safety; Zhang et al. (ZhangDYW23) applied static ownership analysis; Ling et al. (CRustS) and Hong et al. (HongR23; HongR24a; HongR24b; Forcrat) focused on API safety and specific constructs; Larson et al. (Larsen24) proposed pointer derivation graphs to translate unsafe Rust to safer Rust; Wu et al. (WuD25) retyped polymorphic C void pointers into uses of Rust generics. While effective, these methods require significant manual effort and often produce Rust code that is both unsafe and non-idiomatic.
LLM-based Methods leverage LLMs to generate idiomatic Rust code without relying on manually defined rules. Eniser et al. (FLOURINE) introduced FLOURINE, which uses fuzz testing to validate semantic equivalence between C and Rust code. Yang et al. (vert) adapted MSWasm-based testing for multi-language translation to Rust. Luo et al. (IRENE) proposed IRENE, an LLM-based framework that integrates rule-augmented retrieval and structured summarization for function-level translation. Nitin et al. (SpecTra) enhanced LLM translation by integrating static specifications and I/O tests. Xu et al. (XuH25) optimized type migration by leveraging data flow graphs. Farrukh et al. (SafeTrans) developed a multi-agent framework featuring iterative error correction. However, LLM-based translation suffers from lower accuracy due to the lack of parallel C-Rust training data (RustRepoTrans).
Hybrid Methods seek to combine the strengths of both paradigms. For instance, C2SaferRust (C2SaferRust) and PR2 (PR2) employ LLMs to refine rule-based translations, improving safety and idiomaticity while preserving functional equivalence through testing. SACTOR (SACTOR) delegates different tasks to each paradigm, applying C2Rust for data types and an LLM for global variables and functions. Our method extends this direction with a two-stage framework: it first uses linguistic feature mappings (i.e., transformation rules) to guide LLM translation, then applies a compilation-driven repair mechanism combining LLM refinement with static analysis. It achieves an effective balance between translation accuracy and safety in automated C-to-Rust migration.
@@ 6.2. Project-level Code Translation with LLMs
Shiraishi et al. (projectc2rust) pioneered this direction by using project metadata to maintain consistency, though their method prioritizes compilation success over functional correctness. Syzygy (Syzygy) and RustMap (RustMap) employ a strategy of translating both functions and test cases while utilizing feedback mechanisms for repair. Nevertheless, these approaches necessitate complete dependency contexts (occasionally through manual intervention) and exhibit limited scalability when applied to complex industrial codebases. Khatry et al. (CRUST-Bench) and Ou et al. (RustRepoTrans) introduced repository-scale benchmarks and self-repair techniques, yet rely on annotated Rust interfaces rarely available in practice. Hong et al. (typemigrating) addressed this with Tymcrat, a type inference-based method for automating signature translation.

### v4

@@ 4.4. Scalability on Larger Projects (RQ4)
To evaluate the scalability and performance of our method on large-scale codebases, we conduct experiments on 10 larger projects (9,973 - 91,588 LoC) from RepoTransBench (Wang et al., 2024), a benchmark derived from highly-starred GitHub repositories. During module translation, we exclude functions that depend on third-party libraries, as our current implementation primarily supports dependencies on the ISO C standard library. Using DeepSeek-V3 as the backbone model, EvoC2Rust achieves acceptable performance, with an ICompRate of 75.61% and a SafeRate of 96.20%. We do not report line acceptance rates due to the prohibitive cost of manually creating verified Rust references for projects at this scale.
@@ 4.5. Case Study
Figure 2. Translation performance at different project scales.
The figure shows a line chart comparing translation performance across projects with different code lengths. The blue line represents ICompRate, which decreases steadily from 100% for projects under 500 lines of code to about 69% for projects over ten thousand lines. The orange line indicates SafeRate, which starts around 98%, dips for medium-sized projects, and then gradually increases to about 98% for the largest projects. The x-axis categorizes projects into five code-length ranges, from under 500 LOC to over ten thousand LOC. Two y-axes are used to separately display the percentage scales for ICompRate and SafeRate.
Figure 3. Translation time at different project scales.
The figure presents a line chart illustrating translation time per line of code across varying project sizes. The x-axis divides projects into five code-length intervals, ranging from under 500 to over ten thousand lines. The y-axis measures time in seconds per line of code. The trend begins at approximately 0.75 seconds per line for the smallest projects (¡500 LOC), declines to around 0.35 seconds for projects between 501–1000 LOC, peaks at about 0.85 seconds for the 1001–5000 LOC range, and subsequently decreases and stabilizes near 0.55 seconds per line for the largest projects.
@@ 5. Limitations and Threats to Validity
Internal Validity. The main internal threat comes from correctness verification, which currently relies on predefined test cases. Adopting more robust validation techniques like fuzzing (Kadron et al., 2024) and self-debugging (Kang et al., 2025; Shi et al., 2024) could enhance test coverage and uncover subtle semantic errors. For line acceptance rate measurement, we use Claude-generated outputs verified by human experts as ground truth. While this manual validation ensures correctness beyond automated testing, it may not scale to larger datasets. Future work should explore semi-automated verification to balance accuracy and efficiency.
@@ 6.1. C-to-Rust Translation
Rule-based Methods rely on manually designed rules to transform C code into Rust. The C2Rust transpiler (3) exemplifies this approach by using Clang’s AST to generate Rust code, though the output relies heavily on unsafe blocks. Subsequent work has aimed to improve safety and idiomaticity: Emre et al. (Emre et al., 2021, 2023) used compiler feedback for pointer safety; Zhang et al. (Zhang et al., 2023) applied static ownership analysis; Ling et al. (Ling et al., 2022) and Hong et al. (Hong and Ryu, 2023, 2024a, 2024b, 2025a) focused on API safety and specific constructs; Larson et al. (Larsen, 2024) proposed pointer derivation graphs to translate unsafe Rust to safer Rust; Wu et al. (Wu and Demsky, 2025) retyped polymorphic C void pointers into uses of Rust generics. While effective, these methods require significant manual effort and often produce Rust code that is both unsafe and non-idiomatic.
Figure 4. A case study on translating the rb_tree_rotate function.
The figure compares four versions of the rb_tree_rotate function: the original C implementation, a direct prompting translation in Rust, a self-repair version, and the EvoC2Rust result. The top-left panel shows the original C code. The top-right panel displays Rust code generated by direct prompting, which includes unsafe blocks and interface mismatches. The bottom-left panel presents Rust code generated by self-repair, which introduces ownership and move errors. The bottom-right panel depicts the EvoC2Rust translation, which uses consistent pointer casting and produces safe, compilable Rust code. Overall, the figure illustrates how EvoC2Rust successfully avoids the unsafe constructs and ownership pitfalls observed in alternative translation approaches.
LLM-based Methods leverage LLMs to generate idiomatic Rust code without relying on manually defined rules. Eniser et al. (Eniser et al., 2024) introduced FLOURINE, which uses fuzz testing to validate semantic equivalence between C and Rust code. Yang et al. (Yang et al., 2024a) adapted MSWasm-based testing for multi-language translation to Rust. Luo et al. (Luo et al., 2025) proposed IRENE, an LLM-based framework that integrates rule-augmented retrieval and structured summarization for function-level translation. Nitin et al. (Nitin and Ray, 2024) enhanced LLM translation by integrating static specifications and I/O tests. Xu et al. (Xu and Huang, 2025) optimized type migration by leveraging data flow graphs. Farrukh et al. (Farrukh et al., 2025) developed a multi-agent framework featuring iterative error correction. However, LLM-based translation suffers from lower accuracy due to the lack of parallel C-Rust training data (Ou et al., 2024).
Hybrid Methods seek to combine the strengths of both paradigms. For instance, C2SaferRust (Nitin et al., 2025) and PR2 (Gao et al., 2025) employ LLMs to refine rule-based translations, improving safety and idiomaticity while preserving functional equivalence through testing. SACTOR (Zhou et al., 2025) delegates different tasks to each paradigm, applying C2Rust for data types and an LLM for global variables and functions. Our method extends this direction with a two-stage framework: it first uses linguistic feature mappings (i.e., transformation rules) to guide LLM translation, then applies a compilation-driven repair mechanism combining LLM refinement with static analysis. It achieves an effective balance between translation accuracy and safety in automated C-to-Rust migration.
@@ 6.2. Project-level Code Translation with LLMs
Shiraishi et al. (Shiraishi and Shinagawa, 2024) pioneered this direction by using project metadata to maintain consistency, though their method prioritizes compilation success over functional correctness. Syzygy (Shetty et al., 2024) and RustMap (Cai et al., 2025) employ a strategy of translating both functions and test cases while utilizing feedback mechanisms for repair. Nevertheless, these approaches necessitate complete dependency contexts (occasionally through manual intervention) and exhibit limited scalability when applied to complex industrial codebases. Khatry et al. (Khatry et al., 2025) and Ou et al. (Ou et al., 2024) introduced repository-scale benchmarks and self-repair techniques, yet rely on annotated Rust interfaces rarely available in practice. Hong et al. (Hong and Ryu, 2025b) addressed this with Tymcrat, a type inference-based method for automating signature translation.